Discrete Mathematics & Theoretical Computer Science Copyright Transfer

Author(s):

Title of the paper:

Title of the conference:

I hereby transfer to the Discrete Mathematics and Theoretical Computer Science, DMTCS, non profit organization constitued under French law, all the rights concerning the conference paper herein mentionned. I confirm, that my co-authors^{1} and their employer(s)^{1} agreed on this step.

I confirm that the following clauses are part of this agreement:

• The authors keep the right to reuse any portion of the work, without fee, in future works of their own provided that the DMTCS citation are properly included.

• DMTCS guarantees free access to the paper on the internet.

• Copyright falls back to the authors in case the internet service of DMTCS can not be continued.

• The authors, as anybody else, for non-profit use only have the right to copy the electronic version of the paper as it is on the server of DMTCS and to distribute it freely in its electronic form.

Date: Address:

Signature:

Discrete Mathematics and Theoretical Computer Science (DMTCS)

62, rue du Cardinal Mathieu F-54000 Nancy – France

1Please rule out if not applicable.